OpenStudy (anonymous):

factor 24x3y-20x2y2-24xy3

OpenStudy (anonymous):

4xy(6x2 - 5xy - 6y2)

OpenStudy (anonymous):

4xy(6x^2-5xy-6y^2)

OpenStudy (anonymous):

4x3y-20x2y2-24xy3 =4xy(x^2-5xy-6y^2) =4xy(^2-2xy-3xy-6y^2) =4xy(x(x-2y)-3y(x-2y)) =4xy(x-3y)(x-2y)

OpenStudy (anonymous):

sorry my mistake i apologise 4x3y-20x2y2-24xy3 =4xy(x^2-5xy-6y^2) =4xy(x^2-6xy+xy-6y^2) =4xy(x(x-6y)+y(x-6y)) =4xy(x+y)(x-6y)

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

4xy(3x+2Y)(2x-3y)

OpenStudy (anonymous):

= 4xy( 6x2 - 5xy - 6y2 ) = 4xy( 6x2 - 9xy + 4xy - 6y2 ) = 4xy( 3x(2x - 3y) + 2y(2x - 3y)) = 4xy(3x +2y)(2x - 3y)